# A Neural Rewriting System to Solve Algorithmic Problems

### v1

@@ 6 Discussion
In this work, we proposed an original approach to solve formula simplification problems which we dubbed Neural Rewriting System. We described how such problems can be generally solved iteratively by first identifying a leaf formula, then computing its value, and finally replacing the value in the original formula. We show that iteratively decomposing in such a way the solution process and implementing a specialized module for each sub-task, it is possible to achieve higher accuracy on the solution of complex formulas than both a highly-specialized neural architecture trained end-to-end and a general purpose Large Language Model.
Limitations While we consider formula simplification problems with different number and kinds of operands, we also pose several constraints on these formulas. For instance, we limit our study to formulas which have only one final form which can always be reached independently of the substitution applied. Both of these conditions are not satisfied, in general, by algebraic expressions, whose form could be made arbitrarily complex and not further reducible. Future work should thus consider the more general case of formulas with several possible final forms, reachable from different solution paths. This scenario would require learning a decision mechanism to select appropriate rewriting rules, for example by using reinforcement learning. Furthermore, the architectural mechanisms that we employ to generalize on formula simplification problems rely on the exploitation of specific assumptions that apply to the class of problems consider in this study. More general architectural elements and/or training methods should thus be devised to achieve more robust and widely-applicable generalization in neural systems.

### v2

@@ 6 Discussion
In this work, we have considered formula simplification problems, a kind of benchmark used to study the systematic generalization capabilities of neural networks. We proposed a formal description of this class of problems and used it to define a simple general solution algorithm involving three elementary functions. We then implemented a general neural architecture called the Neural Rewriting System, which models such algorithm by adopting a modular design.
We have compared the performance of the Neural Rewriting System to that of a highly specialized neural architecture, the Neural Data Router, and of GPT-4, a powerful general-purpose large language model. We have tested the models on nested arithmetic and algebraic formulas, and formulas from the ListOps dataset. We showed that while both the Neural Data Router and GPT-4 are, to different extents, able to solve simple formulas, both models struggle or fail when dealing with deeply nested formulas. The Neural Rewriting System achieves a considerably greater degree of generalization on much more complex formulas than those it was trained on. We analyzed models’ failures in light of our formalization of formula simplification problems, highlighting the features of these problems that could represent the main obstacles to systematic generalization.
Limitations While we consider formula simplification problems with different number and kinds of operands, we also pose several constraints on these formulas. For instance, we limit our study to formulas with only one final form which can always be reached independently of the substitution applied. Both of these conditions are not satisfied, in general, by algebraic expressions, whose form could be made arbitrarily complex and not further reducible. Future work should thus study the systematic generalization capabilities of neural models in the more general case of formulas with several possible final forms, reachable from different solution paths.
7 Acknowledgements
We thank OpenAI for granting research access to the GPT APIs.